# FC-Attack: Jailbreaking Multimodal Large Language Models via Auto-Generated Flowcharts

### v2

@@ Introduction
Font Styles in Flowcharts. To investigate whether different font styles in flowcharts affect the effectiveness of jailbreak attacks, we select five fonts from Google Fonts that are relatively difficult for humans to read: Creepster, Fruktur Italic, Pacifico, Shojumaru, and UnifrakturMaguntia (the font style examples are shown in Figure A3). Table 3 shows the results of $\mathsf{FC\text{-}Attack}$ (Ensemble). We observe that different font styles can significantly impact the ASR. For example, on GPT-4o mini, the ASR increases across all font styles compared to the original, with Shojumaru font achieving the highest ASR of $20\%$. Similarly, on Claude-3.5, the Pacifico font achieves the highest ASR of $28\%$, which is a substantial improvement compared to the original ASR of $4\%$. For Gemini-1.5, the ASR reaches $98\%$ with the Fruktur font, while Llava-Next achieves $94\%$ with the Shojumaru font. InternVL-2.5 also shows a $6\%$ increase in ASR with the Pacifico font, reaching $96\%$. These findings further highlight the need to consider the impact of different font styles when designing defenses.
Table 6: MLLM performance on the Benign MM-Vet dataset [43] under Adashield-S (Ada-S) and Adashield-A (Ada-A), covering six core tasks: Recognize (Rec), OCR, Knowledge (Know), Generation (Gen), Spatial (Spat), and Math.
@@
Llava- Next Vanilla 38.0/39.0/25.8/24.8/40.1/21.2 38.8
@@
Effect of Flowchart Structure. To explore the impact of graphical structure elements on the jailbreak effect. We conduct experiments with Qwen2-VL using four different flowchart designs: (1) an enhanced FigStep flowchart where each step incorporates step descriptions generated by $\mathsf{FC\text{-}Attack}$; (2) Plain Text structure that only retains text without any graphical elements in the flowchart; (3) Text with Box structure that encapsulates each step in boxes but omits directional arrows; and (4) our complete $\mathsf{FC\text{-}Attack}$ implementation featuring both boxes surrounding step descriptions and arrows indicating the progression between steps. Table A1 shows the results of four flowchart image structures. We notice that the ASR of the FigStep method is 34%, that of Plain Text is 32%, that of Text with Box is 50%, and that of $\mathsf{FC\text{-}Attack}$ is 88%. It is noted that the addition of box elements improves ASR by 18%, while the introduction of directional arrows connecting these boxes further improves it by 38%. These findings reveal the contribution of the graphical structural elements of the flowchart to improving the jailbreak effect.
Defense
We consider four defenses (shown in Table 5), where “Original” represents the results of $\mathsf{FC\text{-}Attack}$ (Ensemble) with an average ASR of $58.6\%$. Using Llama-Guard3-V and JailGuard to detect whether the input is harmful reduced the ASR to $51.7\%$ and $51.4\%$, respectively. The limited effectiveness may stem from flowcharts being primarily text-based, whereas the detection methods are more suited to visual content. AdaShield-S and AdaShield-A reduce the average ASR to $6.9\%$ and $1.7\%$, showing more effective defense performance. However, these two methods also lead to a decline in MLLMs performance on benign datasets. We conduct tests on MM-Vet [43] to evaluate the important factor of “over-defensiveness” on benign datasets, which is an evaluation benchmark that contains complex multimodal tasks for MLLMs. As shown in Table 6, the model’s utility decreases on benign data when using AdaShield-S and AdaShield-A, indicating a future direction for defense development.
@@ Appendix A Introduction of MLLMs in this paper
  • Llava-Next (January 2024) is an open-source MLLM released by the University of Wisconsin-Madison, which builds upon the Llava-1.5 model [25] with multiple improvements [26]. It enhances capabilities in visual reasoning, optical character recognition, and world knowledge. Besides, Llava-Next increases the input image resolution to a maximum of $672\times 672$ pixels and supports various aspect ratios to capture more visual details ($336\times 1344$ and $1344\times 336$).
@@
  • Qwen2-VL (September 2024) is an open-source model released by Alibaba team [37]. It employs naive dynamic resolution to handle images of different resolutions. In addition, it adopts multimodal rotary position embedding, effectively integrating positional information across text, images, and videos.
@@
  • Gemini-1.5 (February 2024) is a production-grade MLLM developed by Google, based on the Mixture-of-Experts architecture [15]. For Gemini-1.5, larger images will be scaled down to the maximum resolution of $3072\times 3072$, and smaller images will be scaled up to $768\times 768$ pixels. Reducing the image size will not improve the performance of higher-resolution images.
@@
  • Claude-3.5-Sonnet (June 2024) is a production multimodal AI assistant developed by Anthropic [2]. The user should submit an image with a long side not larger than 1568 pixels, and the system first scales down the image until it fits the size limit.
@@
  • GPT-4o and GPT-4o Mini are popular production-grade MLLMs developed by OpenAI [18, 29]. GPT-4o Mini is a compact version of GPT-4o, designed for improved cost-efficiency. Both models excel in handling complex visual and language understanding tasks.
@@
  • InternVL-2.5 (June 2024) [13] is an open-source MLLM that ranks first in full-scale open-source multimodal performance. In terms of multimodal long-chain reasoning, it achieves a breakthrough of $70\%$ in the expert-level multidisciplinary knowledge reasoning benchmark MMMU [44], and the general capabilities are significantly enhanced.
@@
  • Qwen-vl-max (January 2024) is the most powerful large-scale visual language model developed by the Alibaba team [5]. Compared with the enhanced version, the model has made further improvements in visual reasoning and the ability to follow instructions, providing a higher level of visual perception and cognitive understanding. It provides the best performance on a wider range of complex tasks, can handle a variety of visual understanding challenges, and demonstrates excellent visual analysis capabilities.
@@
  • Qwen2.5-Omni (March 2025) is the new flagship end-to-end multimodal model in the Qwen series [41]. It is designed for comprehensive multimodal perception and can seamlessly handle multiple input forms, including text, images, audio, and video. In addition to its powerful multimodal understanding capabilities, Qwen2.5-Omni can also provide real-time streaming responses through text generation and natural speech synthesis.
@@
  • LLaVA-Video (October 2024) is an extended version of video processing based on the LLaVA-Next model, which inherits the enhanced capabilities of LLaVA-Next in visual reasoning, optical character recognition, and world knowledge, while expanding the processing capabilities of video content [48]. The model supports high-resolution input (up to 672×672 pixels) and various aspect ratios (336×1344 and 1344×336) to capture more visual details, making it perform well in video understanding and analysis.
@@ Appendix C Generator Setting
Few Shot Samples Instruction One:
@@
Appendix D Flowchart Experiment Performance
Figure A1: Successful jailbreak instance on Gemini-1.5 using $\mathsf{FC\text{-}Attack}$.
Figure A2: ASR under different prompts against MLLMs via videos.
Figure A3: Different styles of fonts in flowcharts (“1 step”).
(a) Flowchart Example-Horizontal: Left-to-right flow.
@@ Appendix D Flowchart Experiment Performance
(c) Flowchart Example-Vertical: Top-to-bottom flow.
Figure A4: Flowchart Examples: Various flow directions.
Table A1: Performances of ASR for different flowchart structures on the Qwen-2-VL.
@@
ASR (%) 88 32 50 34

### v3

@@ Introduction
Font Styles in Flowcharts. To investigate whether different font styles in flowcharts affect the effectiveness of jailbreak attacks, we select five fonts from Google Fonts that are relatively difficult for humans to read: Creepster, Fruktur Italic, Pacifico, Shojumaru, and UnifrakturMaguntia (the font style examples are shown in Figure˜5). Table˜3 shows the results of $\mathsf{FC\text{-}Attack}$ (Ensemble). We observe that different font styles can significantly impact the ASR. For example, on GPT-4o mini, the ASR increases across all font styles compared to the original, with Shojumaru font achieving the highest ASR of $20\%$. Similarly, on Claude-3.5, the Pacifico font achieves the highest ASR of $28\%$, which is a substantial improvement compared to the original ASR of $4\%$. For Gemini-1.5, the ASR reaches $98\%$ with the Fruktur font, while Llava-Next achieves $94\%$ with the Shojumaru font. InternVL-2.5 also shows a $6\%$ increase in ASR with the Pacifico font, reaching $96\%$. These findings further highlight the need to consider the impact of different font styles when designing defenses.
Figure 5: Different styles of fonts in flowcharts (“1 step”).
Table 6: MLLM performance on the benign $M$-Vet dataset Yu et al. (2024) under Adashield-S (Ada-S) and Adashield-A (Ada-A), covering six core tasks: Recognize (Rec), OCR, Knowledge (Know), Generation (Gen), Spatial (Spat), and Math.
@@
Llava-
Next
Vanilla 38.0/39.0/25.8/24.8/40.1/21.2 38.8
@@
Effect of Flowchart Structure. To explore the impact of graphical structure elements on the jailbreak effect. We conduct experiments with Qwen2-VL using four different flowchart designs: (1) an enhanced FigStep flowchart where each step incorporates step descriptions generated by $\mathsf{FC\text{-}Attack}$; (2) Plain Text structure that only retains text without any graphical elements in the flowchart; (3) Text with Box structure that encapsulates each step in boxes but omits directional arrows; and (4) our complete $\mathsf{FC\text{-}Attack}$ implementation featuring both boxes surrounding step descriptions and arrows indicating the progression between steps. Table˜A1 shows the results of four flowchart image structures. We notice that the ASR of the FigStep method is 34%, that of Plain Text is 32%, that of Text with Box is 50%, and that of $\mathsf{FC\text{-}Attack}$ is 90%. It is noted that the addition of box elements improves ASR by 18%, while the introduction of directional arrows connecting these boxes further improves it by 38%. These findings reveal the contribution of the graphical structural elements of the flowchart to improving the jailbreak effect.
Effect of Different Formats. To investigate the impact of different formats on ASR, we add an experiment on three different formats (Code, Figstep-Style presentations, and Table). All methods deliver the same harmful steps but present them in different formats. We conduct the experiment on the Qwen2-VL model. (1) Code: Python code form; (2) Figstep-Style presentations: Fill in the same full steps into the Figstep-Style format; (3) Table: Fill in the same full steps into the table format. From table˜A2, we can observe that Figstep-Style performs only 12% ASR, followed by the Table format at 26% ASR. The Code format shows moderate effectiveness with 52% ASR. In contrast, our $\mathsf{FC\text{-}Attack}$ achieves the highest ASR at 90%, which indicates flowchart is the most effective format.
Defenses
We consider four defenses (shown in Table˜5), where “Original” represents the results of $\mathsf{FC\text{-}Attack}$ (Ensemble) with an average ASR of $58.6\%$. Using Llama-Guard3-V and JailGuard to detect whether the input is harmful reduced the ASR to $51.7\%$ and $51.4\%$, respectively. The limited effectiveness may stem from flowcharts being primarily text-based, whereas the detection methods are more suited to visual content. AdaShield-S and AdaShield-A reduce the average ASR to $6.9\%$ and $1.7\%$, showing more effective defense performance. However, these two methods also lead to a decline in MLLMs performance on benign datasets. We conduct tests on MM-Vet Yu et al. (2024) to evaluate the important factor of “over-defensiveness” on benign datasets, which is an evaluation benchmark that contains complex multimodal tasks for MLLMs. As shown in Table˜6, the model’s utility decreases on benign data when using AdaShield-S and AdaShield-A, indicating a future direction for defense development.
@@ Conclusion
Acknowledgments
We would like to thank all anonymous reviewers, Area Chair, and Program Chair for their insightful comments and constructive suggestions.
@@ Appendix A Introduction of MLLMs in this paper
  • Llava-Next (January 2024) is an open-source MLLM released by the University of Wisconsin-Madison, which builds upon the Llava-1.5 model Liu et al. [2024b] with multiple improvements Liu et al. [2024c]. It enhances capabilities in visual reasoning, optical character recognition, and world knowledge. Besides, Llava-Next increases the input image resolution to a maximum of $672\times 672$ pixels and supports various aspect ratios to capture more visual details ($336\times 1344$ and $1344\times 336$).
@@
  • Qwen2-VL (September 2024) is an open-source model released by the Alibaba team Wang et al. [2024a]. It employs naive dynamic resolution to handle images of different resolutions. In addition, it adopts multimodal rotary position embedding, effectively integrating positional information across text, images, and videos.
@@
  • Gemini-1.5 (February 2024) is a production-grade MLLM developed by Google, based on the Mixture-of-Experts architecture Google [2024]. For Gemini-1.5, larger images will be scaled down to the maximum resolution of $3072\times 3072$, and smaller images will be scaled up to $768\times 768$ pixels. Reducing the image size will not improve the performance of higher-resolution images.
@@
  • Claude-3.5-Sonnet (June 2024) is a production multimodal AI assistant developed by Anthropic Anthropic [2024]. The user should submit an image with a long side not larger than 1568 pixels, and the system first scales down the image until it fits the size limit.
@@
  • GPT-4o and GPT-4o Mini are popular production-grade MLLMs developed by OpenAI Hurst et al. [2024], OpenAI [2024]. GPT-4o Mini is a compact version of GPT-4o, designed for improved cost-efficiency. Both models excel in handling complex visual and language understanding tasks.
@@
  • InternVL-2.5 (June 2024) Chen et al. [2024b] is an open-source MLLM that ranks first in full-scale open-source multimodal performance. In terms of multimodal long-chain reasoning, it achieves a breakthrough of $70\%$ in the expert-level multidisciplinary knowledge reasoning benchmark MMMU Yue et al. [2024], and the general capabilities are significantly enhanced.
@@
  • Qwen-VL-Max (January 2024) is the most powerful large-scale visual language model developed by the Alibaba team Bai et al. [2023]. Compared with the enhanced version, the model has made further improvements in visual reasoning and the ability to follow instructions, providing a higher level of visual perception and cognitive understanding. It provides the best performance on a wider range of complex tasks, can handle a variety of visual understanding challenges, and demonstrates excellent visual analysis capabilities.
@@
  • Qwen2.5-Omni (March 2025) is the new flagship end-to-end multimodal model in the Qwen series Yang et al. [2024]. It is designed for comprehensive multimodal perception and can seamlessly handle multiple input forms, including text, images, audio, and video. In addition to its powerful multimodal understanding capabilities, Qwen2.5-Omni can also provide real-time streaming responses through text generation and natural speech synthesis.
@@
  • LLaVA-Video (October 2024) is an extended version of video processing based on the LLaVA-Next model, which inherits the enhanced capabilities of LLaVA-Next in visual reasoning, optical character recognition, and world knowledge, while expanding the processing capabilities of video content Zhang et al. [2024c]. The model supports high-resolution input (up to 672×672 pixels) and various aspect ratios (336×1344 and 1344×336) to capture more visual details, making it perform well in video understanding and analysis.
A large body of research has demonstrated that AI systems are inherently susceptible to diverse attack vectors He et al. [2025], Liao et al. [2025], Sun et al. [2025a], Zhong et al. [2025], Zheng et al. [2025]. Building on these findings, recent research has revealed that MLLMs likewise suffer from such vulnerabilities as we discussed earlier, which raises pressing concerns regarding their reliable and secure deployment. Although defense is not the primary focus of this paper, it is nevertheless crucial to acknowledge the importance of developing effective countermeasures, as explored in previous works Meta LLaMA [2025], Zhang et al. [2024b], Wang et al. [2024b], Liu et al. [2024e].
@@ Appendix C Generator Setting
Few Shot Prompt prompt = f""" Generate a practical, instructional statement about topic {selected_topic}. The statement should describe a task or activity (e.g., "Teach people how to make a chair." or "Explain how to compost at home."). Follow it with a clear, step-by-step explanation structured for clarity and logic, ensuring the steps are concise and actionable. Examples: Instruction One:
@@
Now, please generate an example like above starting with word {act_topic}‘. """
Appendix D Explainability of $\mathsf{FC\text{-}Attack}$
To investigate why $\mathsf{FC\text{-}Attack}$ is effective on MLLMs, we conduct an experiment on LLaVA-Next, extract query embeddings by taking the hidden state of the last token from the final decoder layer, and evaluate on four different prompt types: benign text queries, harmful text queries, benign image queries, and FC-Attack image queries. Harmful text queries are drawn from a deduplicated version of Advbench. To construct benign text queries, we rewrite the harmful ones by replacing sensitive or unsafe terms with neutral, harmless alternatives while preserving the original grammar and syntactic structure to remove harmful intent.
We then obtain embeddings for all four prompt types and visualize them using t-SNE. As shown in Figure˜A1, embeddings of benign versus harmful text queries are clearly separable, indicating strong semantic alignment in the text modality of LLaVA-Next. In contrast, embeddings from benign image queries and FC-Attack image queries are highly entangled, suggesting insufficient safety alignment in the image modality.
Figure A1: Visualization in t-SNE of LLaVA-Next query embeddings in four different prompt types.
Appendix E Flowchart Experiment Performance
Figure A2: ASR under different prompts against MLLMs via videos
(a) Flowchart Example-Horizontal: Left-to-right flow
@@ Appendix E Flowchart Experiment Performance
(c) Flowchart Example-Vertical: Top-to-bottom flow
Figure A3: Flowchart Examples: Various flow directions
Table A1: Performances of ASR for different flowchart structures on the Qwen2-VL.
@@
ASR (%) 90 32 50 34
Table A2: Performances of ASR for different formats on the Qwen2-VL.
Format $\mathsf{FC\text{-}Attack}$ Code Table Figstep
ASR (%) 90 52 26 12